# Exact Sequence Interpolation with Transformers

### v2

@@ 1.5. Parameter Complexity of Transformers versus ResNets
where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is realized by a (possibly deep) ResNet. This architecture acts independently on each token, which poses a fundamental limitation for solving ˜1.2 under ˜1.3. To see this, consider two input sequences that differ only in their last token,
@@
so one can never simultaneously have $\operatorname{\mathrm{T}}_{\mathrm{FF}}(X^{1})=Y^{1}$ and $\operatorname{\mathrm{T}}_{\mathrm{FF}}(X^{2})=Y^{2}$. Thus, ˜1.2 cannot be solved by feed-forward only transformers in general.
We next discuss a second approach, which ignores sequence structure altogether. Setting $n=\max_{j\in[N]}n^{j}$, $m=\max_{j\in[N]}m^{j}$, and padding shorter sequences with zeros, each input $X^{j}$ can be regarded as a vector in $\mathbb{R}^{dn}$. Similarly, each output $Y^{j}$ can be viewed as a vector in $\mathbb{R}^{dm}$. In this setting, ResNets operate on $\mathbb{R}^{dn}$, so their parameter count necessarily scales with $n$. For example, the ResNets constructed in [29] require $\mathcal{O}(Ndn)$ parameters to interpolate the dataset, while our transformers require $\mathcal{O}(d\sum_{j=1}^{N}m^{j})$ parameters. This comparison yields two regimes:
@@
The efficiency gain of transformers for shorter output sequences stems from the clustering effect of self-attention (see, e.g., [11, 13, 2, 18, 5]). Tokens within a sequence are carefully collapsed to be equal to the output sequence in the set sense. ResNets, in contrast, rely only on geometric correlations in the input, so parameter reductions can be established only probabilistically [35].
The key advantage of transformers is that they combine the token-level processing of ResNets with self-attention driven for separation and clustering. This structure enables exact interpolation for sequences of arbitrary input length, with parameter complexity depending only on the total output length. The tradeoff, which we do not address here, is the quadratic computational cost of self-attention with respect to the input sequence length [31].
1.6. Organization of the Paper
The remainder of the paper is organized as follows. In Section˜2 we examine the implications of our exact interpolation results for regularized training. We then introduce the transformer architecture we analyze in Section˜3. Section˜4 collects auxiliary technical results that will be used to prove Theorem˜1.4 in Section˜5 and Theorem˜1.6 in Section˜6. Finally, Section˜7 discusses perspectives and open problems.
2. Implications for Transformer Regularized Training
@@
(2.1) $\min_{\theta\in\mathbb{R}^{P}}\left\{F_{\varepsilon}(\theta)=\frac{1}{N}\sum_{j=1}^{N}f(\operatorname{\mathrm{T}}_{\theta}(X^{j}),Y^{j})+\varepsilon\kappa(\theta)\right\},$
where $f:\mathcal{X}\times\mathcal{X}\to[0,\infty)$ is a nonnegative lower semicontinuous function such that $f(Y,Y^{\prime})=0$ if and only if $Y=Y^{\prime}$, $\varepsilon>0$ is a regularization parameter, and the function $\kappa:\mathbb{R}^{P}\to[0,\infty)$ is lower semicontinuous, has compact sublevel sets and is coercive (i.e. $\lim_{\|\theta\|\to\infty}\kappa(\theta)=\infty$). The assumptions on $\kappa$ are needed to ensure the existence of a minimizer of (2.1), denoted as $\theta_{\varepsilon}^{\star}$.
@@
(2.2) $\min_{\theta\in\mathbb{R}^{P}}\kappa(\theta)\quad\text{s.t. }\operatorname{\mathrm{T}}_{\theta}(X^{j})=Y^{j}\quad\forall j\in[N],$
as established by the following standard result in Tikhonov regularization theory (see, e.g., [10]).
@@
    There exists $\theta^{\star}_{0}$ solving ˜$2.$
@@
    The limit of any converging subsequence $\{\theta^{\star}_{\varepsilon}\}_{\varepsilon>0}$ solves ˜$2.$
@@
Several remarks on this result are in order. First, the statement has been formulated for softmax transformers since their continuous dependence on $\theta$, when composed with a suitable lower semicontinuous function $f$, guarantees the lower semicontinuity of $\theta\mapsto F_{\varepsilon}(\theta)$. Consequently, a minimizer of ˜2.1 exists in this setting. In contrast, the discontinuity inherent to hardmax transformers prevents such a guarantee.
@@
Second, Theorem˜1.6 establishes the existence of a parameter $\theta_{0}$ that exactly interpolates the dataset of $N$ sequences, provided the softmax transformer has width $d^{\prime}\geq 4$ and $L\geq 2\sum_{j=1}^{N}m^{j}+3N$ blocks. As a result, the conclusions of ˜2.1 apply directly to this model class.
@@
Finally, although the content of ˜2.1 is relatively standard, it carries practical implications for the training dynamics. In particular, if optimization converges to a global minimizer of the regularized objective ˜2.1, then the training loss scales linearly with the regularization parameter $\varepsilon$ as $\varepsilon\to 0$. Conversely, deviations from this behavior indicate either that the optimization procedure converges only to a local minimizer, or that exact interpolation is not possible with the chosen model architecture. We illustrate this phenomenon for transformer models with a synthetic example [^2] Code used to reproduce the example can be found at: https://github.com/DCN-FAU-AvH/exactInterpTF.
@@
Figure 1. In the left panel, the training loss (log scale) for Example˜$2.$ The red dotted line corresponds to the Tikhonov threshold $\varepsilon\|\theta_{\text{exact}}\|_{2}^{2}$. In the right panel, the minimum of the training loss for different choices of the regularization parameter $\varepsilon$.
@@
Since each output sequence consist of a single point, we define the loss function as
@@
which is easily verified to be continuous (and hence lower semicontinuous). We take $\kappa(\theta)=\|\theta\|_{2}^{2}$ as the $\ell^{2}$ norm in the vectorized parameter space $\mathbb{R}^{P}$. We train the model with Adam [20] for $5\cdot 10^{4}$ optimization steps (epochs), and regularization parameter $\varepsilon=10^{-4}$.
To evaluate whether the optimization is approaching a global minimizer, we compare the training loss to the theoretical upper bound $\varepsilon\|\theta_{\text{exact}}\|_{2}^{2}$ deduced from ˜2.1(ii) for our choice of $\kappa$. As shown in the left panel of Figure˜1, the regularized loss falls below this bound only past approximately $10^{4}$ epochs. We repeat the experiment for various values of the regularization parameter $\varepsilon\in\{10^{-1},10^{-2},10^{-3}\}$, recording the minimal loss achieved by the optimizer for each. The results, displayed in the right panel of Figure˜1, exhibit an approximately linear decay consistent with the theoretical prediction in ˜2.1. We conclude that the training process is converging either to a global minimizer, or to a local minimizer with the same $\varepsilon$-scaling for the range of $\varepsilon$ we considered.
3. The Transformer Architecture
In this section, we describe the transformer architecture used to solve ˜1.2. The transformer is defined as a composition of transformer blocks, in which a feed-forward (FF) layer is followed by a self-attention (SA) layer. Both layers may or may not include a residual connection. We define these layers in detail next.
3.1. Feed-Forward Layers
@@
We stress that allowing for a tunable residual parameter $\eta$, rather than fixing $\eta=1$, is crucial to obtain the exact interpolation results in Theorem˜1.4 and Theorem˜1.6. In particular, it allows the composition of feed-forward layers to behave as multi-layer perceptrons (MLPs) and ResNets, whose interpolation capabilities have been studied extensively [15, 29, 3].
3.2. Self-Attention Layers
Self-attention layers are defined as functions $\mathbb{SA}:\mathcal{X}\to\mathcal{X}$ parametrized by matrices $V,A\in\mathbb{R}^{d\times d}$ and $\rho\in\mathbb{R}$ as follows. The $i$-th component of the sequence
@@ 3.2. Self-Attention Layers
(3.2) $\pi_{i\ell}^{\tau}(X,A)=\frac{\exp{\left(\frac{1}{\tau}\langle Ax_{i},x_{\ell}\rangle\right)}}{\sum_{k=1}^{\operatorname{len}(X)}\exp{\left(\frac{1}{\tau}\langle Ax_{i},x_{k}\rangle\right)}}.$
By taking $\tau\to 0$ for fixed $X$ and $A$, one obtains the so-called hardmax self-attention formulation [9]. We denote it by $\pi=\pi^{0}$, where
@@
Figure 2. Geometric interpretation of ˜3.3 for $i=1$ with $A=I$. Tokens $x_{2}$ and $x_{3}$ have the largest orthogonal projection onto $Ax_{1}=x_{1}$, so $\mathcal{C}_{i}(X,A)=\{2,3\}$.
Although rarely used in practice, we use the hardmax self-attention formulation in our proofs due to its simple geometric interpretation: a token $x_{i}$ is influenced by the tokens with the largest orthogonal projection onto the direction of $Ax_{i}$ (cf. Figure˜2). The smooth softmax self-attention formulation is usually preferred in applications because it allows for training with standard gradient-based algorithms.
3.3. The Transformer
We are now ready to combine FF and SA layers into a transformer. Fix a depth $L\in\mathbb{N}$, FF layers $\{\mathbb{FF}^{k}\}_{k\in[L]}$ with parameters $\{\eta^{k},W^{k},U^{k},b^{k}\}_{k\in[L]}$ and constant width $d^{\prime}\in\mathbb{N}$, and SA layers $\{\mathbb{SA}^{k}\}_{k\in[L]}$ with parameters $\{\rho^{k},V^{k},A^{k}\}_{k\in[L]}$. A transformer is a map $\operatorname{\mathrm{T}}:\mathcal{X}\to\mathcal{X}$ defined by
(3.4) $\operatorname{\mathrm{T}}=\left(\mathbb{SA}^{L}\circ\mathbb{FF}^{L}\right)\circ\dots\circ\left(\mathbb{SA}^{1}\circ\mathbb{FF}^{1}\right).$
For each $k\in[L]$, we call the function $\mathbb{TB}^{k}=\mathbb{SA}^{k}\circ\mathbb{FF}^{k}$ a transformer block and say that $\operatorname{\mathrm{T}}$ in ˜3.4 has $L$ blocks (see Figure˜3).

### v3

@@ 1.5 Parameter complexity of transformers versus ResNets
where $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is realized by a (possibly deep) ResNet. This architecture acts independently on each token, which poses a fundamental limitation for solving 1.2 under 1.5. To see this, consider two input sequences that differ only in their last token,
@@
so one can never simultaneously have $\operatorname{\mathrm{T}}_{\mathrm{FF}}(X^{1})=Y^{1}$ and $\operatorname{\mathrm{T}}_{\mathrm{FF}}(X^{2})=Y^{2}$. Thus, 1.2 cannot be solved by feed-forward only transformers in general.
We next discuss a second approach, which ignores sequence structure altogether. Setting $n=\max_{j\in[N]}n^{j}$, $m=\max_{j\in[N]}m^{j}$, and padding shorter sequences with zeros, each input $X^{j}$ can be regarded as a vector in $\mathbb{R}^{dn}$. Similarly, each output $Y^{j}$ can be viewed as a vector in $\mathbb{R}^{dm}$. In this setting, ResNets operate on $\mathbb{R}^{dn}$, so their parameter count necessarily scales with $n$. For example, the ResNets constructed in [domenec2023NODES] require $\mathcal{O}(Ndn)$ parameters to interpolate the dataset, while our transformers require $\mathcal{O}(d\sum_{j=1}^{N}m^{j})$ parameters. This comparison yields two regimes:
@@
The efficiency gain of transformers for shorter output sequences stems from the clustering effect of self-attention (see, e.g., [geshkovski2023emergence, geshkovski2025mathematical, alcalde2025clustering, karagodin2024clustering, burger2025analysis]). Tokens within a sequence are carefully collapsed to be equal to the output sequence in the set sense. ResNets, in contrast, rely only on geometric correlations in the input, so parameter reductions can be established only probabilistically [alvarez2025clusterClassification].
The key advantage of transformers is that they combine the token-level processing of ResNets with self-attention driven for separation and clustering. This structure enables exact interpolation for sequences of arbitrary input length, with parameter complexity depending only on the total output length. The tradeoff, which we do not address here, is the quadratic computational cost of self-attention with respect to the input sequence length [vaswaniAttentionAllYou2017].
1.6 Organization of the paper
The remainder of the paper is organized as follows. In Section 2 we examine the implications of our exact interpolation results for regularized training. We then introduce the transformer architecture we analyze in Section 3. Section 4 collects auxiliary technical results that will be used to prove Theorem 1.6 in Section 5 and Theorem 1.8 in Section 6. Finally, Section 7 discusses perspectives and open problems.
2 Implications for transformer regularized training
@@
$\min_{\theta\in\mathbb{R}^{P}}\left\{F_{\varepsilon}(\theta)=\frac{1}{N}\sum_{j=1}^{N}f(\operatorname{\mathrm{T}}_{\theta}(X^{j}),Y^{j})+\varepsilon\kappa(\theta)\right\},$ (2.1)
where $f:\mathcal{X}\times\mathcal{X}\to[0,\infty)$ is a nonnegative lower semicontinuous function such that $f(Y,Y^{\prime})=0$ if and only if $Y=Y^{\prime}$, $\varepsilon>0$ is a regularization parameter, and the function $\kappa:\mathbb{R}^{P}\to[0,\infty)$ is lower semicontinuous and coercive (i.e. $\lim_{\|\theta\|\to\infty}\kappa(\theta)=\infty$). These assumptions ensure the minimum in (2.1) is attained by a minimizer $\theta_{\varepsilon}^{\star}$.
@@
$\min_{\theta\in\mathbb{R}^{P}}\kappa(\theta)\quad\text{s.t. }\operatorname{\mathrm{T}}_{\theta}(X^{j})=Y^{j}\quad\forall j\in[N],$ (2.2)
as established by the following standard result in Tikhonov regularization theory (see, e.g., [Engl1996Regularization]). We include the proof to make this paper self-contained.
@@
    There exists $\theta^{\star}_{0}$ solving $2.$
@@
    The limit of any converging subsequence $\{\theta^{\star}_{\varepsilon}\}_{\varepsilon>0}$ solves $2.$
Proof.
To prove i), note that any minimizing sequence $\{\theta_{n}\}$ for problem 2.2 remains in the sublevel set $\{\theta:\kappa(\theta)\leq\kappa(\theta_{0})\}$, which is compact because $\kappa$ is coercive and lower semicontinuous. Pass to a convergent subsequence (not relabeled) and denote its limit by $\theta_{0}^{*}$. This point is feasible for 2.2 because the map $\theta\mapsto\operatorname{\mathrm{T}}_{\theta}(X^{j})$ is continuous for every fixed $X^{j}$. It also satisfies $\kappa(\theta_{0}^{*})\leq\liminf_{n}(\theta_{n})$ because $\kappa$ is lower semicontinuous. Then, since $\{\theta_{n}\}$ is a minimizing sequence, $\theta_{0}^{*}$ must be an optimal solution of 2.2.
Next, we prove ii). Since $\theta^{\star}_{0}$ is admissible but not necessarily optimal for 2.1, we obtain
$F_{\varepsilon}(\theta^{\star}_{\varepsilon})\leq F_{\varepsilon}(\theta^{\star}_{0})=\frac{1}{N}\sum_{j=1}^{N}f(\mathrm{T}_{\theta_{0}^{\star}}(X^{j}),Y^{j})+\varepsilon\kappa(\theta^{\star}_{0})=\varepsilon\kappa(\theta^{\star}_{0}).$
The second equality is true because $\mathrm{T}_{\theta_{0}^{\star}}(X^{j})=Y^{j}$ and $f(Y^{j},Y^{j})=0$.
Finally, we prove iii) and iv). Since $f$ is nonnegative, we have that $\varepsilon\kappa(\theta^{\star}_{\varepsilon})\leq F_{\varepsilon}(\theta^{\star}_{\varepsilon})$, which by part ii) implies that $\kappa(\theta^{\star}_{\varepsilon})\leq\kappa(\theta^{\star}_{0})$. Thus, $\{\theta^{\star}_{\varepsilon}\}_{\varepsilon>0}$ is uniformly bounded and has a converging subsequence $\{\theta_{\varepsilon_{j}}\}$. Let $\theta^{*}$ be its limit. Since $\theta\mapsto\mathrm{T}_{\theta}(X^{j})$ is continuous and $f$ is lower semicontinuous, the function
$\theta\mapsto\sum_{i=1}^{N}f(\mathrm{T}_{\theta}(X^{j}),Y^{j})$
is lower semicontinuous. We then find that
$0\leq\sum_{i=1}^{N}f(\mathrm{T}_{\theta^{*}}(X^{j}),Y^{j})\leq\liminf_{\varepsilon_{j}\to 0}F_{\varepsilon_{j}}(\theta^{\star}_{\varepsilon_{j}})\leq\liminf_{\varepsilon_{j}\to 0}\varepsilon_{j}\|\theta^{\star}_{0}\|=0.$
We deduce that $f(\mathrm{T}_{\theta^{*}}(X^{j}),Y^{j})=0$ for all $j\in[N]$, so $\mathrm{T}_{\theta^{*}}(X^{j})=Y^{j}$ for all $j\in[N]$. Thus, $\theta^{*}$ is feasible for problem 2.2. It is also optimal because $\kappa$ is lower semicontinuous and $\kappa(\theta^{*}_{\varepsilon_{j}})\leq\kappa(\theta_{0}^{*})$, so $\kappa(\theta^{*})\leq\liminf\kappa(\theta^{\star}_{\varepsilon_{j}})\leq\kappa(\theta^{\star}_{0})$. The proof is complete. ∎
@@
Several remarks on 2.1 are in order. First, the statement has been formulated for softmax transformers since their continuous dependence on $\theta$, when composed with a suitable lower semicontinuous function $f$, guarantees the lower semicontinuity of $\theta\mapsto F_{\varepsilon}(\theta)$. Consequently, a minimizer of 2.1 exists in this setting. In contrast, the discontinuity inherent to hardmax transformers prevents such a guarantee.
@@
Second, Theorem 1.8 establishes the existence of a parameter $\theta_{0}$ that exactly interpolates the dataset of $N$ sequences, provided the softmax transformer has width $d^{\prime}\geq 4$ and $L\geq 2\sum_{j=1}^{N}m^{j}+3N$ blocks. As a result, the conclusions of 2.1 apply directly to this model class.
@@
Finally, although the content of 2.1 is relatively standard, it carries practical implications for the training dynamics. In particular, if optimization converges to a global minimizer of the regularized objective 2.1, then the training loss scales at worst linearly with the regularization parameter $\varepsilon$ as $\varepsilon\to 0$. Conversely, deviations from this behavior indicate either that the optimization procedure converges only to a local minimizer, or that exact interpolation is not possible with the chosen model architecture. We illustrate this phenomenon for transformer models with a synthetic example [^1] Code used to reproduce the example can be found at: https://github.com/DCN-FAU-AvH/exactInterpTF.
@@
Figure 1: In the left panel, the training loss (log scale) for Example 2.2. The red dotted line corresponds to the Tikhonov threshold $\varepsilon\|\theta_{\text{exact}}\|_{2}^{2}$. In the right panel, the minimum of the training loss for different choices of the regularization parameter $\varepsilon$.
@@
Since each output sequence consists of a single point, we define the loss function as
@@
which is easily verified to be continuous (and hence lower semicontinuous). We take $\kappa(\theta)=\|\theta\|_{2}^{2}$ as the $\ell^{2}$ norm in the vectorized parameter space $\mathbb{R}^{P}$. We train the model with Adam [kingma2014adam] for $5\cdot 10^{4}$ optimization steps (epochs), and regularization parameter $\varepsilon=10^{-4}$.
To evaluate whether the optimization is approaching a global minimizer, we compare the training loss to the theoretical upper bound $\varepsilon\|\theta_{\text{exact}}\|_{2}^{2}$ deduced from 2.1(ii) for our choice of $\kappa$. As shown in the left panel of Figure 1, the regularized loss falls below this bound only past approximately $10^{4}$ epochs. We repeat the experiment for various values of the regularization parameter $\varepsilon\in\{10^{-1},10^{-2},10^{-3}\}$, recording the minimal loss achieved by the optimizer for each. The results, displayed in the right panel of Figure 1, exhibit an approximately linear decay consistent with the theoretical prediction in 2.1. We conclude that the training process is converging either to a global minimizer, or to a local minimizer with the same $\varepsilon$-scaling for the range of $\varepsilon$ we considered.
3 The transformer architecture
In this section, we describe the transformer architecture used to solve 1.2. The transformer is defined as a composition of transformer blocks, in which a feed-forward (FF) layer is followed by a self-attention (SA) layer. Both layers may or may not include a residual connection. We define these layers in detail next.
3.1 Feed-forward layers
@@
We stress that allowing for a tunable residual parameter $\eta$, rather than fixing $\eta=1$, is crucial to obtain the exact interpolation results in Theorem 1.6 and Theorem 1.8. In particular, it allows the composition of feed-forward layers to behave as multi-layer perceptrons (MLPs) and ResNets, whose interpolation capabilities have been studied extensively [hernandez2024deep, domenec2023NODES, alvarez2024interplay].
3.2 Self-attention layers
Let $\mathcal{V},\mathcal{A}$ be two subspaces of $\mathbb{R}^{d\times d}$. Self-attention layers are defined as functions $\mathbb{SA}:\mathcal{X}\to\mathcal{X}$ parametrized by a scalar $\rho\in\mathbb{R}$, a matrix $V\in\mathcal{V}$ and a matrix $A\in\mathcal{A}$ as follows. The $i$-th component of the sequence
@@ 3.2 Self-attention layers
$\pi_{i\ell}^{\tau}(X,A)=\frac{\exp{\left(\frac{1}{\tau}\langle Ax_{i},x_{\ell}\rangle\right)}}{\sum_{k=1}^{\operatorname{len}(X)}\exp{\left(\frac{1}{\tau}\langle Ax_{i},x_{k}\rangle\right)}}.$ (3.2)
Transformer implementations typically parameterize the attention scores using query and key matrices $Q,K\in\mathbb{R}^{h\times d}$, $h\leq d$ [vaswaniAttentionAllYou2017], with similarities given by $\langle Qx_{i},Kx_{\ell}\rangle$. Our formulation absorbs these projections into a single matrix $A=K^{\top}Q$ and reduces precisely to this setting when $\mathcal{A}$ is the subspace of $d\times d$ matrices with rank $h$.
By taking $\tau\to 0$ for fixed $X$ and $A$, one obtains the so-called hardmax self-attention formulation [elfadelSoftmax1993]. We denote it by $\pi=\pi^{0}$, where
@@
Figure 2: Geometric interpretation of 3.3 for $i=1$ with $A=I$. Tokens $x_{2}$ and $x_{3}$ have the largest orthogonal projection onto $Ax_{1}=x_{1}$, so $\mathcal{C}_{i}(X,A)=\{2,3\}$.
Although rarely used in practice, we use the hardmax self-attention formulation in our proofs due to its simple geometric interpretation: a token $x_{i}$ is influenced by the tokens with the largest orthogonal projection onto the direction of $Ax_{i}$ (cf. Figure 2). The smooth softmax self-attention formulation is usually preferred in applications because it allows for training with standard gradient-based algorithms.
3.3 The transformer
We are now ready to combine FF and SA layers into a transformer. Fix a depth $L\in\mathbb{N}$ and a width $d^{\prime}\in\mathbb{N}$. Given FF layers $\{\mathbb{FF}^{k}\}_{k\in[L]}$ with parameters $\{\eta^{k},W^{k},U^{k},b^{k}\}_{k\in[L]}$ and SA layers $\{\mathbb{SA}^{k}\}_{k\in[L]}$ with parameters $\{\rho^{k},V^{k},A^{k}\}_{k\in[L]}$, a transformer is a map $\operatorname{\mathrm{T}}:\mathcal{X}\to\mathcal{X}$ defined by
$\operatorname{\mathrm{T}}=\left(\mathbb{SA}^{L}\circ\mathbb{FF}^{L}\right)\circ\dots\circ\left(\mathbb{SA}^{1}\circ\mathbb{FF}^{1}\right).$ (3.4)
For each $k\in[L]$, we call the function $\mathbb{TB}^{k}=\mathbb{SA}^{k}\circ\mathbb{FF}^{k}$ a transformer block and say that $\operatorname{\mathrm{T}}$ in 3.4 has $L$ blocks (see Figure 3). Note that, from a mathematical perspective, the order of composition of SA and FF layers is irrelevant because one can always choose the parameters of these layers to make them act like the identity map. In particular, all of our results remain true if the transformer blocks are defined as $\mathbb{TB}^{k}=\mathbb{FF}^{k}\circ\mathbb{SA}^{k}$, as in the work introducing the transformer [vaswaniAttentionAllYou2017].